show how to solve 2x/3 - 3 = 8-3x

Answer by edjones(8007)

2x/3 - 3 = 8-3x
2x-9=24-9x multiply each side of the equation by 3.
11x=33 Add 9x+9 to each side.
x=3
.
Ed